How many moles of NaOCl must be added to 150 mL of 0.025 M HOCl to obtain a buffer solution with a pH = 7.50? Ka= 2.8×10-8 for HOCl.

The answer is 3.3x10-3
Answer #1

PKa = -logka

        = -log2.8*10-8

       = 7.5528

no of moles of HOCl = Molarity * voluem in L

                             = 0.025*0.15 = 0.00375 moles

PH   = PKa + log[NaOCl]/[HOCl

7.5    = 7.5528 + log[NaOCl]/0.00375

log[NaOCl]/0.00375 = 7.5-7.5528

log[NaOCl]/0.00375 = -0.0528

[NaOCl]/0.00375        = 10-0.0528

[NaOCl]/0.00375      = 0.8855

[NaOCl ]              = 0.88*0.00375

no of moles of NaOCl = 0.0033moles   = 3.3*10-3